OpenStudy (anonymous):

Which of the following ordered pairs is a solution to the inequality y -2x + 10? (0, -4) (1, 13) (6, -1) (3, 8)

OpenStudy (anonymous):

There is no inequality in your question. \[\le, \ge, <, or >? \]

OpenStudy (anonymous):

less than or equal to

OpenStudy (anonymous):

so \[y \le -2x+10\] All you have to do is plug in the points and see if the equation makes sense. For (0,-4) you end up with \[-4\le10\] which is true. Next (1,13) you get \[13 \le 8\] Which is false, because 13 is larger than 8. Try the rest on your own.